# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 5.2 Convergence of tree statistics
An important class of these integral functionals, which in fact motivated our choice of the form of convergence of the random measures $\mathcal{L}^{N,n}$, are the internal and external branch lengths of the coalescent associated with the SFS. Formally, the $r$th total branch length $\tau^{n,r}$ of a coalescent $\chi^{n}\in\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$ of a sample of size $n$ is defined by
@@
That is, $\tau^{n,r}$ is the sum of the lengths of the edges in the coalescent ancestral to $r$ samples. Note that $\tau^{n,r}$ is not of the form of $\varphi$ described above, so it requires an additional argument to show that these functionals do indeed converge. This is done via a truncation argument.
@@
Suppose that, as $N\to\infty$, that Assumptions 3.3, 3.4, and 3.6 hold. Then, for any finite $s\in\mathbb{N}$,
@@
By Remark 4.8, the occupation measures $\mu_{\bar{\chi}^{N,n}}$ converge vaguely to $\mu_{\chi^{n}}$, so for any bounded continuous cutoff $\psi_{T,\varepsilon}\in C_{c}(\mathbb{R}_{+})$ with $\int_{0}^{\infty}|\psi_{T,\varepsilon}(s)-\mathbf{1}_{[0,T]}(s)|ds\leq s$,
$\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\bar{\chi}^{N,n}(t))\,dt\,\mid\,\mathcal{A}_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\chi^{n}(t))\,dt\,\mid\,Pi\right].$
@@ 5.2.1 The site-frequency spectrum under nearly complete selfing
Here we illustrate this using simulations of the (robust) conditional limit described in Theorem 5.2, namely, the coalescing random walk on the ancestral recombination graph (ARG). In this limit, the ARG representing all the information of the pedigree relevant to the sample is generated by an EFC in which every distinct pair of ancestral lineages coalesce with rate $2$ and every single lineage split at rate $\lambda$. After an exponentially distributed waiting time with rate parameter $n(\lambda+n-1)$, one of these $\binom{n}{2}+n$ events is chosen in proportion to its rate. If two lineages coalesce, the number of ancestral lineages increases by one. If a lineage splits, the number of ancestral lineages increases by one. All events and their times are recorded, and the process stops the first time there is just a single ancestral lineage.
@@
The gene genealogy at a locus is generated by tracing ancestral lineages backward in time through the graph, as in the presentation of Figure 4. When a lineage encounters a split, it follows one or the other ancestral line with equal probability, $1/2$. When two lineages meet at a coalescent event in the graph, they necessarily coalesce because $\alpha_{N}\to 1$ under limited outcrossing. Unlinked loci have independent gene genealogies, conditional on the graph. Figure 5 displays simulation results for the SFS for six different values of $\lambda$ ranging from $1000$ to 0. Each panel shows the SFS for five independently generated ancestral graphs assuming a sample of size $n=20$. Given each graph, gene genealogies of $10^{5}$ unlinked loci were simulated and their values of $\tau^{n,r}$ were recorded. Figure 5 plots the averages $\overline{\tau^{n,r}}$ of these branch lengths over the $10^{5}$ gene genealogies, normalized to sum to one, that is divided by $\sum_{r}\overline{\tau^{n,r}}$.
In Figure 5(a), which has $\lambda=1000$, the SFS for every graph is not noticeably different than under the Kingman coalescent (not shown). We note that this is also the expectation for a fixed $\alpha_{N}=\alpha$ ND (97); NWF (25). A small effect of the graph (or pedigree) can be seen even when $\lambda=100$ in Figure 5(b). At the other extreme, when $\lambda=0$ as in Figure 5(f), the entire pedigree of the population is reduced to a single Kingman coalescent tree, and the gene genealogies of every locus have this exact same tree. Here, the SFS reflects the random outcomes, of branch lengths and numbers of descendants of each branch, of an effectively single-locus coalescent process. Something quite similar is seen with $\lambda=0.1$ in Figure 5(e). The middle plots, Figure 5(c) and Figure 5(d), with $\lambda=10$ and $\lambda=1$, occupy a transition zone between the deterministic case ($\lambda\to\infty$) where all graphs have the same SFS and the highly stochastic case ($\lambda\to 0$) where each random graph has a distinctly different SFS.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 5: Expected site-frequency spectra under the limited-outcrossing model for 55 independently generated ancestral graphs, corresponding to 55 pedigrees, for each of six different values of $\lambda\in\{1000,100,10,1,0.1,0\}$ in decreasing order from panel (a) to (f). SFS on the vertical axis means the relative expected number of polymorphic sites where the mutant allele is found in each count $r\in\{1,2,3,\ldots,19\}$ in a sample of size $n=20$. These were estimated by simulating gene genealogies of $10^{5}$ unlinked loci on each graph/pedigree.
@@ 5.3 Diploid Sargasyan-Wakeley model
where we noted that $2\frac{K}{N}\frac{N-K}{N-1}$ and $\frac{K(K-1)}{N(N-1)}$ are respectively the probability that exactly one (of the two distinct individuals) is a child and the probability that both individuals are children.
@@
Suppose, as $N\to\infty$, that Assumptions 3.3 and 3.4 hold, and that $c_{N}^{-1}\mu_{N}$ converges vaguely on $0,1]\times\mathbb{N}$ to a finite measure $\mu$. Define the measure $\Xi^{\prime}(\mu)$ on $\Delta$ by
$\int_{(0,1]\times\mathbb{N}}\frac{x^{2}}{m}\,\delta_{\big(\frac{x}{m},\ldots,\frac{x}{m},0,0,\ldots\big)}\,d\mu(x,m),$
where the dirac mass has $m$ copies of $\frac{x}{m}$ in a row and then all zeroes. Then, where $\Pi$ is an EFC with $c_{k}=2(1-\Xi^{\prime}(\mu)(\Delta))$ and $\mu_{\nu_{\rm Coag}}=2\Xi^{\prime}(\mu)$, we have that $\mathcal{L}^{N,n}$ converges weakly as a random measure to $\mathcal{L}_{\Pi}^{n}$.
@@
To prove the claim we proceed by characterizing the ordered offspring distribution $\Phi_{N}$ on $\Delta$ from Lemma Appendix A.1. That is, let $V_{(1)}$, $V_{(2)}$, $\ldots$, $V_{(N)}$ denote the ordered genetic contribution of individuals between time-steps. $\Phi_{N}$ is defined to be the law of the infinite tuple in $\Delta$
$\left(\frac{V_{(1)}}{2N},\frac{V_{(2)}}{2N},\ldots,\frac{V_{(N)}}{2N},0,0,\ldots\right).$
Lemma Appendix A.1 would imply that Assumption 3.6 holds, and that $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent where $\Xi=2\left(\Xi^{\prime}(\mu)+(1-\Xi^{\prime}(\mu))\delta_{\mathbf{0}}\right)$. An application of Theorem 4.7 then gives the result. It suffices, therefore, to show that the conditions of Lemma Appendix A.1 hold.
We will see that, conditional on $(x,m)$ in $(0,1]\times\mathbb{Z}_{+}$, the difference between $\frac{V_{(1)}}{2N}$ and $\frac{V_{(m)}}{2N}$ is negligible as $N\to\infty$. We demonstrate this with the slightly simpler case of putting balls into boxes. The variance of putting $\lfloor xN\rfloor$ balls into $m$ buckets is of order $\frac{\lfloor xN\rfloor}{m}$. When renormalizing by $2N$, the variance is of order $O(\frac{1}{N})$. This shows that the difference between the proportion of the $N$ balls when $\lfloor xN\rfloor$ of the balls are dropped into $P$ boxes for the box with the most and the box with the least vanish as $N\to\infty$. Further, this proportion converges to $\frac{x}{m}$. This is not the exact process for $V_{(i)}$, though the argument carries over in much the same way regardless of which $\alpha$ to which $\alpha_{N}\to\alpha$. However, as $\alpha_{N}\to 1$ by Assumption 3.4 we have that the comparison can be made directly to balls in boxes simply by disregarding the proportion of balls that are the result of outcrossing.
By the above reasoning and the assumption that $c_{N}^{-1}\mu_{N}\to\mu$, Lemma Appendix A.1 yields that Assumption 3.6 holds for $Q_{n}$ the infinitesimal generator of an $n$-$\Xi$-coalescent governed by $\Xi=2(\Xi^{\prime}(\mu)+(1-\Xi^{\prime}(\mu))\delta_{\mathbf{0}})$. This gives the claim.
@@ 5.4 Mixed demographies
Many models of interest involve mixed demographic behavior, such as Kingman-like behavior with rare extreme events, as in BBE (13); DFBW (24). Suppose that we have two different parent-child distributions $\pi_{N}^{(1)}$ and $\pi_{N}^{(2)}$ on $N]\times[N]$ for the Sargasyan-Wakeley model of Example 2.2. Let $c_{N}^{(1)}$ and $c_{N}^{(2)}$ denote the one-step transition probabilities for two sample lineages in distinct individuals coalescing in a single time-step. Similarly, we let $d_{N}^{(1)}$ and $d_{N}^{(2)}$ denote the outcrossing probabilities for a single time-step for each of the two demographies. For simplicity, we assume that the selfing probabilities $\alpha_{N}^{(1)}$ and $\alpha_{N}^{(2)}$ for the two demographic histories are identical, i.e. $\alpha_{N}^{(1)}=\alpha_{N}^{(2)}=\alpha_{N}$. For any fixed $\rho>0$ we define the mixed parent-child distribution $\pi_{N}$ by
$\pi_{N}=(1-\rho c_{N}^{(1)})\pi_{N}^{(1)}+\rho c_{N}^{(1)}\pi_{N}^{(2)}.$
@@
Suppose that, as $N\to\infty$, $c_{N}^{(1)}\to 0$, $\rho c_{N}^{(2)}\to\rho^{\prime}<\infty$, that $\pi_{N}^{(i)}$ satisfies Assumption 3.6 and the measure on $\Delta$ governing the coalescences of the demography of $\pi_{N}^{(i)}$ is $\Xi^{(i)}$ for each $i\in\{1,2\}$, that $\left(c_{N}^{(1)}\right)^{-1}d_{N}^{(1)}\to\lambda<\infty$, and that $d_{N}^{(2)}\to 0$. Then the random measure
$\mathbb{P}\left(\bar{\chi}^{N,n}\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm mixed}\right)$
converges weakly in distribution to $\mathcal{L}_{\Pi}^{n}$, where $\Pi$ is an EFC with characteristics $c_{k}=\Xi^{(1)}(\mathbf{0})+\rho^{\prime}\Xi^{(2)}(\mathbf{0})$, $\nu_{Coag}=\Xi^{(1)}+\rho^{\prime}\Xi^{(2)}-c_{k}\delta_{\mathbf{0}}$, $c_{e}=\lambda$, and $\nu_{Disl}=0$.
@@
As the two demographies satisfy Assumption 3.6 so to does $\pi_{N}$. If $Q_{n}^{(i)}$ is the generator of the demography $\pi_{N}^{(i)}$ then we have that the generator of the mixed demography $Q_{n}^{\rm mixed}$ is
$Q_{n}^{\rm mixed}=Q_{n}^{(1)}+\rho^{\prime}Q_{n}^{(2)}.$
@@
$\Xi^{\rm mixed}=\Xi^{(1)}+\rho^{\prime}\Xi^{(2)}.$
As $d_{N}^{(1)}(c_{N}^{(1)})^{-1}\to\lambda<\infty$ and $d_{N}^{(2)}\to 0$ the outcrossing probability of the mixed demography $d_{N}^{\rm mixed}$ satisfies $d_{N}^{\rm mixed}\left(c_{N}^{(1)}\right)^{-1}\to\lambda$. The result then follows directly from Theorem 4.7. ∎
@@
where $\psi$ is a fixed number in $[0,1]$. That is, with probability $1-\frac{\rho}{N^{2}}$ a single time-step consists of a single reproductive event, a single child born to any one of the $N$ potential parents with probability $\alpha_{N}$ or to any of the $\binom{N}{2}$ pairs of parents with probability $1-\alpha_{N}$; that is, it behaves as a Moran model as in NWF (25). With probability $\frac{\rho}{N^{2}}$, two parents reproduce to form $\lfloor\psi N\rfloor$ of the individuals in the nest time-step (going forward in time.) Let $\mathcal{A}_{N}^{\rm BBE}$ denote the pedigree associated to $\pi_{N}^{\rm BBE}$.
@@
$\mathbb{P}\left(\tilde{\chi}^{N,n}\left(\lfloor t\left(c_{N}^{(1)}\right)^{-1}\rfloor\right)\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm BBE}\right)$
converges weakly in distribution as a random measure to $\mathcal{L}_{\Pi}^{n}$, where $\Pi$ is the EFC of Example 4.3.
@@
and we see that $\rho c_{N}^{(2)}\to\rho\frac{\psi^{2}}{2}$
@@
Suppose now that $\eta$ consists of $b$ blocks $C_{,}C_{2},...,C_{b}$ of sizes $k_{1},k_{2},\ldots,k_{b}$, at most two of which are non-singletons. Where $s$ is the number of singletons, we write that $\eta$ is of type $b;k_{1},\ldots,k_{r};s$, as in Sch (00). We only need to consider the cases now where $s\geq b-2$. Observe that the probability that any of the $l$ individuals containing a sample lineage is one of the two parents is $o(1)$, which will be negligible for $\pi_{N}^{(2)}$. We proceed now assuming that $k_{1}\geq 2$ and that $k_{1}\geq k_{2}\geq\ldots\geq k_{b}$ without loss of generality.
For any $j$ of the $l$ sample lineages, the probability that they are the child of one of the two parents is $\psi^{j}(1-\psi)^{l-j}+o(1)$. If $k_{2}\geq 2$, then there is only one way for $\eta$ to appear, by $k_{1}$ sample lineages falling into one parent, $k_{2}$ in the other, and the remaining $l-k_{1}-k_{2}$ remaining sample lineages not being children. Therefore
@@
Finally, we calculate $q^{l,2}_{\xi_{0}^{l}\xi_{0}^{l}}$. There can be zero, one, or two of the $l$ sample lineages that enter the two parents. Zero lineages entering the two parents occurs with probability $(1-\psi)^{l}+o(1)$. There are $\binom{l}{1}$ ways for $1$ sample lineage to be in one of the two parents, and so this occurs with probability $\binom{l}{1}\psi(1-\psi)^{l-1}+o(1)$. There are $\binom{l}{2}$ ways for $2$ sample lineages to be in the two parents, and in this case they do not coalesce together with probability $\frac{1}{2}$. Therefore

### v2

@@ 5.2 Convergence of tree statistics
An important class of these integral functionals, which in fact motivated our choice of the form of convergence of the random measures $\mathcal{L}^{N,n}$, are the internal and external branch lengths of the coalescent associated with the site-frequency spectrum (SFS). Formally, the $r$th total branch length $\tau^{n,r}$ of a coalescent $\chi^{n}\in\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)$ of a sample of size $n$ is defined by
@@
That is, $\tau^{n,r}$ is the sum of the lengths of the edges in the coalescent ancestral to $r$ samples; see Figure 5 for an illustration.
Figure 5: A gene genealogy for $n=10$ samples. Here, $\tau^{10,3}$ is the sum of the lengths of the two thick, blue edges. Although we do not consider mutation in this paper, note that any mutations on these branches will be inherited by, and thus present on, exactly 33 gene copies in the sample.
Note that $\tau^{n,r}$ is not of the form of $\varphi$ described above, so it requires an additional argument to show that these functionals do indeed converge. This is done via a truncation argument.
@@
Suppose that, as $N\to\infty$, that Assumptions 3.4, 3.5, and 3.7 hold. Then, for any finite $s\in\mathbb{N}$,
@@
By Remark 4.9, the occupation measures $\mu_{\bar{\chi}^{N,n}}$ converge vaguely to $\mu_{\chi^{n}}$, so for any cutoff function $\psi_{T,\varepsilon}\in C_{c}(\mathbb{R}_{+})$ with $\int_{0}^{\infty}|\psi_{T,\varepsilon}(s)-\mathbf{1}_{[0,T]}(s)|ds\leq s$,
$\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\bar{\chi}^{N,n}(t))\,dt\,\mid\,\mathcal{A}_{N}\right]\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{d}}}}\,\mathbb{E}\left[\int_{0}^{\infty}\psi_{T,\varepsilon}(t)g_{r}(\chi^{n}(t))\,dt\,\mid\,\Pi\right].$
@@ 5.2.1 The site-frequency spectrum under nearly complete selfing
Here we illustrate this using simulations of the (robust) conditional limit described in Theorem 5.2, namely, the coalescing random walk on the ancestral graph with binary mergers. In this limit, the ancestral graph representing all the information of the pedigree relevant to the sample is generated by an EFC in which every distinct pair of ancestral lineages coalesce with rate $2$ and every single lineage splits at rate $\lambda$. After an exponentially distributed waiting time with rate parameter $n(\lambda+n-1)$, one of these $\binom{n}{2}+n$ events is chosen in proportion to its rate. If two lineages coalesce, the number of ancestral lineages decreases by one. If a lineage splits, the number of ancestral lineages increases by one. All events and their times are recorded, and the process stops the first time there is just a single ancestral lineage.
@@
The gene genealogy at a locus is generated by tracing ancestral lineages backward in time through the graph, as in the presentation of Figure 4. When a lineage encounters a split, it follows one or the other ancestral line with equal probability, $1/2$. When two lineages meet at a coalescent event in the graph, they necessarily coalesce. Unlinked loci have independent gene genealogies, conditional on the graph. Figure 6 displays simulation results for the SFS for six different values of $\lambda$ ranging from $1000$ to 0. Each panel shows the SFS for five independently generated ancestral graphs assuming a sample of size $n=20$. Given each graph, gene genealogies of $10^{5}$ unlinked loci were simulated and their values of $\tau^{n,r}$ were recorded. Figure 6 plots the averages $\overline{\tau^{n,r}}$ of these branch lengths over the $10^{5}$ gene genealogies, normalized to sum to one, that is divided by $\sum_{r}\overline{\tau^{n,r}}$.
In Figure 6A, which has $\lambda=1000$, the SFS for every graph is not noticeably different than under the Kingman coalescent (not shown). We note that this is also the expectation for a fixed $\alpha_{N}=\alpha$ ND (97); NWF (25). A small effect of the graph (or pedigree) can be seen even when $\lambda=100$ in Figure 6B. At the other extreme, when $\lambda=0$ as in Figure 6F, the entire pedigree of the population is reduced to a single Kingman coalescent tree, and the gene genealogies of every locus have this exact same tree. Here, the SFS reflects the random outcomes, of branch lengths and numbers of descendants of each branch, of an effectively single-locus coalescent process. Something quite similar is seen with $\lambda=0.1$ in Figure 6E. The middle plots, Figure 6C and Figure 6D, with $\lambda=10$ and $\lambda=1$, occupy a transition zone between the deterministic case ($\lambda\to\infty$) where all graphs have the same SFS and the highly stochastic case ($\lambda\to 0$) where each random graph has a distinctly different SFS.
Figure 6: Expected site-frequency spectra under the limited-outcrossing model for 55 independently generated ancestral graphs, corresponding to 55 pedigrees, for each of six different values of $\lambda\in\{1000,100,10,1,0.1,0\}$ in decreasing order from panel A to F. SFS on the vertical axis means the relative expected number of polymorphic sites where the mutant allele is found in each count $r\in\{1,2,3,\ldots,19\}$ in a sample of size $n=20$. These were estimated by simulating gene genealogies of $10^{5}$ unlinked loci on each graph/pedigree.
@@ 5.3 Diploid Sargasyan-Wakeley model
where we noted that $2\frac{K}{N}\frac{N-K}{N-1}$ and $\frac{K(K-1)}{N(N-1)}$ are respectively the probability that exactly one (of the two distinct individuals) is an offspring and the probability that both individuals are offspring.
@@
Suppose, as $N\to\infty$, that Assumptions 3.4 and 3.5 hold, and that $\frac{1}{2c_{N}}\mu_{N}$ converges vaguely on $0,1]\times\mathbb{N}$ to a finite measure $\mu$ for which $\int\frac{x^{2}}{m}d\mu(x,m)\leq 1$. Define the sub-probability measure $\Theta(\mu)$ on $\Delta\setminus\{\bf{0}\}$ by
$\Theta(\mu):=\int_{(0,1]\times\mathbb{N}}\frac{x^{2}}{m}\,\delta_{\big(\frac{x}{m},\ldots,\frac{x}{m},0,0,\ldots\big)}\,d\mu(x,m)\qquad\text{on }\Delta\setminus\{\bf{0}\},$
where the dirac mass has $m$ copies of $\frac{x}{m}$ in a row and then all zeroes. Then, where $\Pi$ is an EFC with $c_{k}=2\big(1-\Theta(\mu)(\Delta\setminus\{\bf{0}\})\big)$ and $\mu_{\nu_{\rm Coag}}=2\Theta(\mu)$, we have that $\mathcal{L}^{N,n}$ converges weakly in law to $\mathcal{L}_{\Pi}^{n}$.
@@
To prove the claim we proceed by characterizing the ordered offspring distribution $\Phi_{N}$ on $\Delta$ from Lemma Appendix B.2. That is, let $\tilde{V}_{(1)}$, $\tilde{V}_{(2)}$, $\ldots$, $\tilde{V}_{(N)}$ denote the ordered genetic contribution of individuals between time-steps as in Lemma Appendix B.2. $\Phi_{N}$ is defined to be the law of the infinite tuple
$\mathfrak{V}_{N}=\left(\frac{\tilde{V}_{(1)}}{2N},\frac{\tilde{V}_{(2)}}{2N},\ldots,\frac{\tilde{V}_{(N)}}{2N},0,0,\ldots\right)\in\Delta.$
We claim that as $N\to\infty$,
$\frac{1}{2c_{N}}\Phi_{N}(dx)\to\frac{1}{\left\langle x,x\right\rangle}\Theta(\mu)(dx)$
vaguely on $\Delta\setminus\{\mathbf{0}\}$. Lemma Appendix B.2 would imply that Assumption 3.7 holds, and that $Q_{n}$ is the infinitesimal generator of an $n$-$\Xi$-coalescent with
$\Xi=2\Theta(\mu)+2\Big(1-\Theta(\mu)(\Delta\setminus\{\bf{0}\})\Big)\,\delta_{\mathbf{0}},$
giving the desired formula for $c_{k}$ and $\mu_{\nu_{\rm Coag}}$.
An application of Theorem 4.7 then gives the result. It suffices, therefore, to show that the conditions of Lemma Appendix B.2 hold.
By definition of vague convergence it suffices to show that, for any $f=g(x_{1},\ldots,x_{M})$ a function of the first $M$ entries of an element $x$ in $\Delta$ with $f(\mathbf{0})=0$ that
$\frac{1}{2c_{N}}\Phi_{N}(f)\to\int\frac{f(z)}{\left\langle z,z\right\rangle}\Theta(\mu)(dz).$
Let $\iota:(0,1]\times\mathbb{N}\to\Delta$ denote the embedding
$(x,m)\mapsto\left(\frac{x}{m},\ldots,\frac{x}{m},0,\ldots\right),$
where $(x,m)$ is mapped to the element of $\Delta$ consisting of $m$ copies of $\frac{x}{m}$ and then infinitely many zeroes. With this notation we have that
$\int\frac{f(z)}{\left\langle z,z\right\rangle}\Theta(\mu)(dz)=\int\frac{f\circ\iota(x,m)}{\frac{x^{2}}{m}}d\Theta(\mu)(x,m)=\int f\circ\iota(x,m)d\mu(x,m)=\mu(f\circ\iota).$
The convergence will follow from a concentration of measure result, which says that
$\mathbb{P}\left(\left|\left|\mathfrak{V}_{N}-\iota(dx,m)\right|\right|\geq\epsilon\,\mid\,\pi_{N}\in[x,x+dx]\times\{m\})\right)\leq\frac{1}{\epsilon}O\left(\frac{1}{N}\right).$ (5.3)
That is $\mathfrak{V}_{N}$, conditional on the number of parents and offspring $\pi_{N}$ being in the rectangle $dx\times\{m\}$, concentrates around $\iota(dx,m)$. This implies the main result as it implies, by continuity of $f$ and compactness of $\Delta$, that
$\frac{1}{2c_{N}}\Phi_{N}(f)=\frac{1}{2c_{N}}\mu_{N}(f\circ\iota)+o(1)\to\mu(f\circ\iota),$
where the last arrow follows from the assumption that $\frac{1}{2c_{N}}\mu_{N}$ converges vaguely to $\mu$ on $(0,1]\times\mathbb{N}$.
We now justify (5.3). Conditional on $\pi_{N}\in[x,x+dx]\times\{m\}$, exactly $\lfloor dx\,N\rfloor$ diploid offspring are produced by the $m$ potential parents. Ignoring outcrossing events for the moment, the vector of parental contributions $(W_{1},\ldots,W_{m})$ is multinomially distributed with parameters $\lfloor dx\,N\rfloor$ and cell probabilities $(\frac{1}{m},\ldots,\frac{1}{m})$. In particular,
$\mathbb{E}\left[W_{i}\right]=\frac{dx\,N}{m}$
and
$Var\left(W_{i}\right)=\frac{dx\,N}{m}\left(1-\frac{1}{m}\right).$
When we renormalize by $2N$, we obtain
$Var\left(\frac{W_{i}}{2N}\right)=O\left(\frac{x}{mN}\right).$
Chebyshev’s inequality for each fixed $i\leq m$ and a union bound therefore yield (5.3). This union bound works because $f$ depends only on finitely many components of each element of $\Delta$.
Finally, we need to explain why this argument applies when we allow for outcrossing. Because of Assumption 3.5, the proportion of offspring arising from outcrossing events is $o(1)$. These contribute a vanishingly small proportion of the genetic contributions of each of the parents.
@@ 5.4 Mixed demographies
Many models of interest involve mixed demographic behavior, such as Kingman-like behavior with rare extreme events, as in BBE (13); DFBW (24). Suppose that we have two different parent-offspring distributions $\pi_{N}^{(1)}$ and $\pi_{N}^{(2)}$ on $N]\times[N]$ for the Sargasyan-Wakeley model of Example 2.2.
Let $c_{N}^{(1)}$ and $c_{N}^{(2)}$ denote the one-step transition probabilities for two sample lineages in distinct individuals coalescing in a single time-step. Similarly, we let $d_{N}^{(1)}$ and $d_{N}^{(2)}$ denote the outcrossing probabilities for a single time-step for each of the two demographies. For simplicity, we assume that the selfing probabilities $\alpha_{N}^{(1)}$ and $\alpha_{N}^{(2)}$ for the two demographic histories are identical, i.e. $\alpha_{N}^{(1)}=\alpha_{N}^{(2)}=\alpha_{N}$. For any fixed $\rho>0$ we define the mixed parent-offspring distribution $\pi_{N}$ by
$\pi_{N}=(1-\rho c_{N}^{(1)})\pi_{N}^{(1)}+\rho c_{N}^{(1)}\pi_{N}^{(2)}.$ (5.4)
@@
Suppose that $c_{N}^{(1)}\to 0$, $\rho c_{N}^{(2)}\to\rho^{\prime}<\infty$, $\left(c_{N}^{(1)}\right)^{-1}d_{N}^{(1)}\to\lambda<\infty$, and $d_{N}^{(2)}\to 0$ as $N\to\infty$. Suppose also that, for each $i\in\{1,2\}$, the Sargasyan-Wakeley model (Example 2.2) under $\pi_{N}^{(i)}$ satisfies Assumption 3.7 and the measure on $\Delta$ governing the limiting annealed coalescence is $\Xi^{(i)}$. Then Assumption 3.7 holds for the Sargasyan-Wakeley model under $\pi_{N}$ defined by (5.4), and
$\mathbb{P}\left(\tilde{\chi}\left(\lfloor tc_{N}^{(1)}(1+\rho^{\prime})\rfloor\right)\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm mixed}\right)\,{\mathrel{\mathop{\kern 0.0pt\longrightarrow}\limits^{{wd\;}}}}\,\mathcal{L}_{\Pi}^{n}\quad\text{in}\quad\mathcal{M}_{1}(\mathcal{D}\left(\mathbb{R}_{+},\mathcal{E}_{n}\right)),$
where $\Pi$ is an EFC with characteristics
$c_{k}=$ $\frac{1}{1+\rho^{\prime}}\Xi^{(1)}(\mathbf{0})+\frac{\rho^{\prime}}{1+\rho^{\prime}}\Xi^{(2)}(\mathbf{0}),$ (5.5)
$\nu_{Coag}=$ $\frac{1}{1+\rho^{\prime}}\Xi^{(1)}+\frac{\rho^{\prime}}{1+\rho^{\prime}}\Xi^{(2)}-c_{k}\delta_{\mathbf{0}},$ (5.6)
$c_{e}=$ $\frac{\lambda}{1+\rho^{\prime}}\quad\text{and}\quad\nu_{Disl}=0.$ (5.7)
@@
As the two demographies satisfy Assumption 3.7, so does $\pi_{N}$. If $Q_{n}^{(i)}$ is the generator of the demography $\pi_{N}^{(i)}$ then we have that the generator of the mixed demography $Q_{n}^{\rm mixed}$ is
$Q_{n}^{\rm mixed}=\frac{1}{1+\rho^{\prime}}Q_{n}^{(1)}+\frac{\rho^{\prime}}{1+\rho^{\prime}}Q_{n}^{(2)}+o(1),$
as
$c_{N}=(1-\rho c_{N}^{(1)})c_{N}^{(1)}+\rho c_{N}^{(1)}c_{N}^{(2)}=c_{N}^{(1)}(1+\rho^{\prime}+o(1)).$
@@
$\Xi^{\rm mixed}=\frac{1}{1+\rho^{\prime}}\Xi^{(1)}+\frac{\rho^{\prime}}{1+\rho^{\prime}}\Xi^{(2)}.$
As $d_{N}^{(1)}(c_{N}^{(1)})^{-1}\to\lambda<\infty$ and $d_{N}^{(2)}\to 0$ the outcrossing probability of the mixed demography $d_{N}^{\rm mixed}=(1-\rho c_{N}^{(1)})d_{N}^{(1)}+\rho c_{N}^{(1)}d_{N}^{(2)}$ satisfies $d_{N}^{\rm mixed}\left(c_{N}^{(1)}(1+\rho^{\prime})\right)^{-1}\to\frac{\lambda}{1+\rho^{\prime}}$. The result then follows directly from Theorem 4.7. ∎
@@
where $\psi$ is a fixed number in $[0,1]$. That is, with probability $1-\frac{\rho}{N^{2}}$ a single time-step consists of a single reproductive event, a single offspring born to any one of the $N$ potential parents with probability $\alpha_{N}$ or to any of the $\binom{N}{2}$ pairs of parents with probability $1-\alpha_{N}$; that is, it behaves as a Moran model as in NWF (25). With probability $\frac{\rho}{N^{2}}$, two parents reproduce to form $\lfloor\psi N\rfloor$ of the individuals in the nest time-step (going forward in time.) Let $\mathcal{A}_{N}^{\rm BBE}$ denote the pedigree associated to $\pi_{N}^{\rm BBE}$.
@@
$\mathbb{P}\left(\tilde{\chi}^{N,n}\left(\left\lfloor t\left(c_{N}^{(1)}\left(1+\rho\frac{\psi^{2}}{2}\right)\right)^{-1}\right\rfloor\right)\in\cdot\,\mid\,\mathcal{A}_{N}^{\rm BBE}\right)$
converges weakly in law to $\mathcal{L}_{\Pi}^{n}$, where $\Pi$ is the EFC of Example 4.3.
@@
and we see that $\rho c_{N}^{(2)}\to\rho\frac{\psi^{2}}{2}$ (i.e. $\rho^{\prime}=\rho\frac{\psi^{2}}{2}$ in the assumption of Corollary 5.6).
@@
Suppose now that $\eta$ consists of $b$ blocks $C_{1},C_{2},...,C_{b}$ of sizes $k_{1},k_{2},\ldots,k_{b}$, at most two of which are non-singletons. Where $s$ is the number of singletons, we write that $\eta$ is of type $b;k_{1},\ldots,k_{r};s$, as in Sch (00). We only need to consider the cases now where $s\geq b-2$. Observe that the probability that any of the $l$ individuals containing a sample lineage is one of the two parents is $o(1)$ as $N\to\infty$, which will be negligible for $\pi_{N}^{(2)}$. We proceed now assuming that $k_{1}\geq 2$ and that $k_{1}\geq k_{2}\geq\ldots\geq k_{b}$ without loss of generality.
For any $j$ of the $l$ sample lineages, the probability that they are the offspring of one of the two parents is $\psi^{j}(1-\psi)^{l-j}+o(1)$. If $k_{2}\geq 2$, then there is only one way for $\eta$ to appear, by $k_{1}$ sample lineages falling into one parent, $k_{2}$ in the other, and the remaining $l-k_{1}-k_{2}$ remaining sample lineages not being offspring. Therefore
@@
Finally, we calculate $q^{l,2}_{\xi_{0}^{l}\xi_{0}^{l}}$. There can be zero, one, or two of the $l$ sample lineages that enter the two parents. Zero lineages entering the two parents occurs with probability $(1-\psi)^{l}+o(1)$ as $N\to\infty$. There are $\binom{l}{1}$ ways for $1$ sample lineage to be in one of the two parents, and so this occurs with probability $\binom{l}{1}\psi(1-\psi)^{l-1}+o(1)$. There are $\binom{l}{2}$ ways for $2$ sample lineages to be in the two parents, and in this case they do not coalesce together with probability $\frac{1}{2}$. Therefore